Assuming complete evaporation and no air change does the concentration created by the following conditions exceeds the LEL ( lower explosive limit ) . A leak in a pressurized piping system sprays 5 gal. of Ethyl Alcohol in a room 15 ft. x 15 ft. x 8 ft. MW= 46; SG= 0.6; LEL=3.3%.
Answer #1

Volume of room = 15*15* 8= 1800 ft3 since 1ft3= 28.32L

1800 ft3= 1800*28.32 L= 50976 L

5 gallons of alcohol correspond to 5*3.78 L=18.9 L mass of ethyl alcohol =

Volume percent of ethyl alcohol in air = 100*18.9/50976 =0.037 given LEL = 3.3/100=0.033% So more than LEL.

hence conditinos excel LEL.